### v5

@@ 4.2. Extended Transformation
[ReflexivityApp] Γ⊢A ⇑B Γ⊢reflexivity(≡_A) ⇑
Γ reflexivity(≡_B)
@@
Figure 4. The additional rules needed for repairing across setoid equivalences. There are two mutually defined judgements: one to repair environments (top) and one to repair terms (bottom).
We extend this transformation to work for setoid equivalences as well. To do this, we adapt our transformation to handle the changes in how equality works (Figure 4). We handle equivalence and its proofs in three cases: its type, its construction by reflexivity, and its elimination by rewriting.
@@
To define repair, we provide custom equivalence relations for any type we wish to consider as a setoid. This is done by providing three terms:
@@
The next rule deals with constructions of equivalence relations by reflexivity. The proof supplied by the user that $\equiv_{C}$ is an equivalence relation contains a term for reflexivity, with type $\forall$ c : C, $\equiv_{C}$ c c.
We denote this term reflexivity($\equiv_{C}$). If $\equiv_{C}$ is @eq C, this is simply eq_refl C. In either case, each reflexivity($\equiv_{C}$) repairs to reflexivity($\equiv_{D}$) (by ReflexivityApp).
@@ 5. Implementation
We implement this extension to the transformation by extending the Pumpkin Pi Coq plugin (Section 5.1). Consistently with the original Pumpkin Pi, our implementation relies on some user annotations (Section 5.2), and places some other restrictions on the format terms can take (Section 5.3). Our implementation includes novel automation to dispatch proper proofs specific to setoids (Section 5.4).
@@ 5.1. Extending Pumpkin Pi
Pumpkin Pi has various classes of proof repair transformations across type equivalences for which it has specialized automation. We add an additional class, termed setoid lifting, to support our extended transformation. This class mostly reuses the existing transformation, but implements the new rules from Figure 4. In all, our extension adds 1659 lines of code, 510 of which are dedicated to proper proof generation implemented in propergen.ml. The core extensions to the transformation can be found in lift.ml, liftconfig.ml, and liftrules.ml in the supplementary material.
In our extended configuration, the proof that $\equiv_{C}$ is an equivalence relation takes the form of instances of type classes Equivalence $\equiv_{C}$. This makes it possible to use Coq’s setoid automation to implement the oracle that produces proofs of rewrites described in Section 4.1. Coq has a tactic, called setoid_rewrite, which attempts to perform rewriting by an equivalence relation. However, because equivalence relations are not generally Leibniz, we must prove for each function we define that the function is proper, as defined in Section 3, if we wish to rewrite under applications of that function. When we prove this, we instantiate the Proper type class. The setoid_rewrite tactic uses the Proper and Equivalence type class instances to search for proofs of rewrites, and thus we can use it as our oracle.
@@ 5.2. User Annotations
To perform repair, Pumpkin Pi requires that users annotate their proofs explicitly with components of the configuration. These annotations are required to identify parts of the configuration, thereby decoupling the undecidable part of proof repair (configuration inference) from the decidable part (the proof term transformation itself). We inherit this design decision for our extension, for which this applies to both the original configuration and to our extension to equivalences.
For the original configuration, we do not yet implement any new automation for the new class of changes we support to unify arbitrary subterms with applications of DepConstr, DepElim, $\eta$, and $\iota$. Users of repair across setoid equivalences must apply these terms from the configuration directly in order for repair to work. In contrast, for some specific classes of changes implemented in the original Pumpkin Pi work, these annotations can sometimes be automatically inferred, lessening the burden of user annotation.
For the new configuration components corresponding to equivalence relations, we have some inference in simple cases. Specifically, if no relation is provided for a given type, our extension defaults to strict equality. Then, we can automatically infer annotations corresponding to applications of equality @eq C, as well as applications of reflexivity @eq_refl C. The same holds for rewrites that fully apply any of Coq’s equality eliminators. However, this annotation inference cannot deal with unapplied instances of @eq and @eq_refl that may later be specialized to C, or equality eliminators that are not fully applied. Such terms are improperly annotated.
We do not yet have annotation inference for rewrites when the equivalence relation is user-specified and does not correspond to equality, as these setoid rewrites can manifest as long and complex subterms that are difficult to automatically identify. Instead, we supply a custom user annotation START_REWRITE, which the user must apply to their term before applying the rewrite. This constant takes two non-implicit arguments: the proof of equivalence by which the user is rewriting, and the type they are rewriting. We provide a custom tactic rewrite_annotate which the user can use in place of rewrite in proofs, which automatically performs annotation while rewriting.
@@ 5.3. Restrictions
While the original proof repair work had a single dependent eliminator, we have multiple eliminators. One eliminates into the sort Type, but is purely nondependent. The other is dependent, but only eliminates into the sort Prop, and requires that the motive of the eliminator be proven to be proper, considering Prop as a setoid with iff. The types of two of these eliminators for one of our case studies are in Figure 5. We do this because Coq’s setoid automation does not work for a dependently typed notion of an equivalence relation. Thus, we cannot use the setoid automation to perform rewrites on applications of functions we define with a dependently typed eliminator. For our Prop-sorted eliminator, this loss means that users cannot automatically perform rewrites on the proofs of propositions. In Coq, Prop is frequently treated as effectively proof irrelevant, so this loss is more acceptable.
In addition, our use of Coq’s setoid automation is facilitated through the use of Coq’s rewrite tactics. However, these tactics do not allow a rewrite which does not change the goal type, and so neither does our automation for repairing setoid rewrites. Furthermore, setoid rewrites do not allow specifying a motive P. To get around this, a user trying to prove P x -> P y can perform a substitution P[z/y], where z is free in P, and then define Q := fun z => P[z/y]. Then, for another fresh variable w, the user can perform a setoid_rewrite to prove H : $\forall$ (w : B), Q w x -> Q w y, and recover the desired rewrite proof as H y.
@@ 5.4. Automating Proper Proofs
To perform rewrites on a term using Coq’s setoid automation, it is necessary to prove that the functions in that term are proper. Thus, when repairing terms, it is potentially necessary to have proper proofs for every previously repaired function. We implement novel automation that helps prove many functions proper automatically. Our automation must fail in some cases, though, since proving a general function is proper is undecidable. To prove this, for any proposition P : Prop, define f : bool -> Prop, f b = if b then True else P. Generating a proof that f is proper, using bool as a setoid relating true and false, and Prop as a setoid with the relation iff, is equivalent to proving iff P True, which is undecidable.
Presently, our automation proves proper proofs automatically in two practical cases. First, when f is a composition of proper functions, we introduce hypotheses stating that all arguments are equivalent, and then rewrite by these hypotheses. This is the approach taken by Coq’s solve_proper tactic, though we modify it slightly. Coq’s solve_proper fails if any of its inputs do not appear in the body of the function because the rewrite for that argument will fail. To avoid this, we use the try tactical when rewriting. We also try both Coq’s rewrite and setoid_rewrite tactics, while solve_proper only runs setoid_rewrite. While setoid_rewrite supports rewriting under binders in some instances, rewrite succeeds in some instances where setoid_rewrite fails.
Second, suppose that f is an application of the eliminator of some inductive type with a constant motive. Then, if all of the inductive case arguments provided are proper functions, we can prove that the eliminator is a proper function from its base cases to its output. To make this concrete, let C be a setoid with eqC the equivalence relation, fix some P = fun _ => C, and consider:
nat_rect P : $\forall$ (po : C)
(ps : forall (n : nat) (pn : C), C) (n : nat), C
@@
Proper (eq ==> eqC ==> eqC) ps -> $\forall$ (n : nat),
Proper (eqC ==> eqC) (fun po => nat_rect po ps n)
by induction on n. The base case holds by assumption via proper, and the inductive case holds by ps being proper.
This can be done in general for inductive types. Thus, our automation checks if f is an eliminator at the top level, and if so generates and tries to prove such a proper goal. If it succeeds, the automation attempts to show that the inductive cases (in the above example, ps) are proper using the rewriting strategy described above. In our case studies, this solves the proper goals for add_posGZ and add_negsucGZ.
@@
Figure 6. The types of our integer representations. We provide an instance of Equivalence eq_GZ in the case study.
@@ 6. Case Studies
We use our extended version of Pumpkin Pi on three case studies that use quotient type equivalences. First, we conduct repair between two representations of the integers (Section 6.1). Second, we study two common implementations of the queue data structure, and how we can repair from one to the other (Section 6.2). Third, we repair between dense and sparse representations of polynomials with natural number coefficients (Section 6.3). All of the case study examples can be found in more detail in the supplementary material.
@@ 6.1. Adding, Fast and Slow
Our first case study is mathematically motivated. We consider a change in the type representing integers. We repair addition and proofs about addition from one representation to the other. Finally, we recover the repaired proofs for a more efficient version of addition over the repaired type. Our Coq implementation of this case study can be found in grothendieck_int_equivalence_repair_tool.v.
@@
Our first representation, Z, is based on the default implementation of the integers in Cubical Agda: two copies of $\mathbb{N}$ glued back-to-back. We will repair functions and proofs about Z to use a representation that may be more familiar to set theorists: viewing the integers as elements of $\mathbb{N}\times\mathbb{N}/\sim$, where $(x_{1},x_{2})\sim(y_{1},y_{2})\iff x_{1}+y_{2}=x_{2}+y_{1}$. We call the resulting type GZ, and the equivalence relation eq_GZ. The definitions are in Figure 6. Z and GZ are setoid equivalent using the definition from Section 3, by mapping pos n to (n, 0) and negsuc n to (0, S n).
@@
Figure 8. An addition identity whose proof we repaired.
@@
We use appliedDepElimPropZ in our proof of add0LZ, and can repair the term. Presently, an implementation bug only surfacing in this case study forces us to repair the arguments to appliedDepElimPropZ before reconfiguring. Future versions can avoid the need for this workaround by automatically supplying the necessary proper proofs to depElimProp terms.
@@
We adapt our repaired proofs to use the more efficient addition function defined in Figure 9. Consistently with prior work in Pumpkin Pi, to move between slow and fast implementations we take an ad hoc approach. First, we prove that both implementations of addition are pointwise equivalent in our setoid. Then, we rewrite across this equality to get from proofs of repaired theorems defined over slow addition to proofs of corresponding theorems defined over fast addition. For example, we use this methodology to translate the proof of add0LGZ into the proof of theorem:
@@
Figure 10. One list queues and two list queues. We provide an instance of Equivalence eq_queue in the case study.

### v6

@@ 3.2. Extended Transformation
[ReflexivityApp] Γ⊢A ⇑B Γ⊢reflexivity(≡_A) ⇑reflexivity(≡_B)
@@
Figure 5. The additional rules needed for repairing across setoid equivalences. There are two mutually defined judgments: one to repair environments (top) and one to repair terms (bottom).
We now extend this transformation to work for setoid equivalences as well. To do this, we adapt our transformation to handle the changes in how equality works (Figure 5). We handle equivalence and its proofs in three cases: its type, its construction by reflexivity, and its elimination by rewriting.
@@
To define repair, the user must provide custom equivalence relations for any type they wish to consider as a setoid. This is done by providing three terms:
@@
The next rule, ReflexivityApp, deals with constructions of equivalence relations by reflexivity. The proof supplied by the user that $\equiv_{C}$ is an equivalence relation contains a term for reflexivity, with type $\forall$ c : C, $\equiv_{C}$ c c. We denote this term reflexivity($\equiv_{C}$). If $\equiv_{C}$ is @eq C, this is simply eq_refl C. In any case, each reflexivity($\equiv_{C}$) repairs to reflexivity($\equiv_{D}$) (by ReflexivityApp).
@@
Inductive unit :=
| tt.
Theorem rewrite_example :
forall (x : unit), eq x tt -> eq x tt.
Proof.
intros x H.
rewrite H.
reflexivity.
Qed.
fun (x : unit) (H : @eq unit x tt) =>
@eq_ind_r unit tt
(fun x0 : unit => @eq unit x0 tt)
(reflexivity(@eq unit) tt) x H
Inductive unit_two :=
| one
| two.
Definition eq_unit_two (u1 u2 : unit_two) : Prop := True.
fun (x : unit_two)
(H : eq_unit_two x one) =>
@@
(unit_two, one,
(fun x0 : unit_two =>
eq_unit_two x0 one),
(reflexivity(eq_unit_two) one), x, H)]]
: forall x : unit_two,
eq_unit_two x one ->
eq_unit_two x one
Figure 6. A demonstration of repairing a rewrite in a proof. The left side is over the source type, which is simply the type with one element. We provide a sample Coq tactic script which performs a rewrite, and below it is the term produced. The right side shows the repaired term over the target setoid, which is the type with two elements which are equal under the equivalence relation.
[…]
Importantly, this extension to Pumpkin Pi can only affect the trusted computing base of the proof system if the rewrite oracle has trust assumptions. Thus, by enforcing that the implementation of the oracle introduces no such assumptions, using this extension poses no risk of a soundness failure beyond that of the base type theory without any extensions.
@@ 4. Implementation
We implement this extension to the transformation by extending the Pumpkin Pi Coq plugin (Section 4.1). Consistently with the original Pumpkin Pi, our implementation relies on some user annotations (Section 4.2), and places some other restrictions on the format terms can take (Section 4.3). Our implementation includes custom automation to dispatch proper proofs specific to setoids (Section 4.4).
@@ 4.1. Extending Pumpkin Pi
Pumpkin Pi has various classes of proof repair transformations across type equivalences for which it has specialized automation. We add an additional class, termed setoid lifting, to support our extended transformation. This class mostly reuses the existing transformation, but implements the new rules from Figure 5. Once the transformation rules were determined, the implementation was made straightforward thanks to Pumpkin Pi’s extensibility; in all, our extension adds 1659 lines of code, 510 of which are dedicated to proper proof generation implemented in propergen.ml. The core extensions to the transformation can be found in lift.ml, liftconfig.ml, and liftrules.ml in the supplementary material.
In our extended configuration, the proof that $\equiv_{C}$ is an equivalence relation takes the form of instances of type classes Equivalence $\equiv_{C}$. This makes it possible to use Coq’s setoid automation to implement the oracle that produces proofs of rewrites described in Section 3.2 with no axioms or other extensions to the trusted computing base. Coq has a tactic, called setoid_rewrite, which attempts to perform rewriting by an equivalence relation. However, because equivalence relations are not generally Leibniz, we must prove for each function we define that the function is proper, as defined in Section 2, if we wish to rewrite under applications of that function. In Coq, we prove this by creating an instance of the Proper type class. The setoid_rewrite tactic uses the Proper and Equivalence type class instances to search for proofs of rewrites, and thus we can use it as our oracle.
@@ 4.2. User Annotations
To perform repair, Pumpkin Pi requires that users annotate their proofs explicitly with components of the configuration. These annotations are required to identify parts of the configuration, thereby decoupling the undecidable part of proof repair (configuration inference) from the decidable part (the proof term transformation itself). We inherit this requriement for our extension, and thus both the original configuration and to our extension to equivalences must be annotated. However, we have implemented automation into Pumpkin Pi that generates many of the needed annotations for equivalence relations automatically.
Specifically, if no relation is provided for a given type, our extension defaults to strict equality. Then, we can automatically infer annotations corresponding to applications of equality @eq C, as well as applications of reflexivity @eq_refl C. The same holds for rewrites that fully apply any of Coq’s equality eliminators. However, this annotation inference cannot deal with unapplied instances of @eq and @eq_refl that may later be specialized to C, or equality eliminators that are not fully applied. Such terms are considered improperly annotated.
When an equivalence relation is provided for a type, we provide a custom tactic rewrite_annotate which the user can use in place of rewrite in proofs, which automatically performs annotation while rewriting. This inserts a custom annotation term START_REWRITE before applying the rewrite, which our extension looks for to identify and repair rewrites. This constant takes two non-implicit arguments: the proof of equivalence by which the user is rewriting, and the type they are rewriting.
@@ 4.3. Restrictions
While the original proof repair work had a single depElim term on each side of the configuration, we have multiple eliminators. One eliminates into the sort Type, but is purely nondependent. The other is dependent, but only eliminates into the sort Prop, and requires that the motive of the eliminator be proven to be proper, considering Prop as a setoid with iff. The types of two of these eliminators for one of our case studies are in Figure 7. We do this because Coq’s setoid automation does not work for a dependently typed notion of an equivalence relation. Thus, we cannot use the setoid automation to perform rewrites on applications of functions we define with a dependently typed eliminator. For our Prop-sorted eliminator, this loss means that users cannot automatically perform rewrites on the proofs of propositions. In Coq, Prop is frequently treated as effectively proof irrelevant, so this loss is more acceptable.
In addition, our use of Coq’s setoid automation is facilitated through the use of Coq’s rewrite tactics. However, these tactics do not allow a rewrite which does not change the goal type, and so neither does our automation for repairing setoid rewrites. Furthermore, the setoid rewrite automation does not allow specifying a motive P. To perform setoid rewrites with a motive, a user trying to prove P x -> P y can perform a substitution P[z/y], where z is free in P, and then define Q := fun z => P[z/y]. Then, for another fresh variable w, the user can perform a setoid_rewrite to prove H : $\forall$ (w : B), Q w x -> Q w y, and recover the desired rewrite proof as H y. Our extension uses this methodology when repairing rewrites with a motive to setoids.
@@ 4.4. Automating Proper Proofs
To perform rewrites on a term using Coq’s setoid automation, it is necessary to prove that the functions in that term are proper. Thus, when repairing terms, it is potentially necessary to have proper proofs for every previously repaired function. We implement automation that helps prove many functions to be proper automatically. Our automation must fail in some cases, though, since proving a general function is proper is undecidable. To prove this, for any proposition P : Prop, define f : bool -> Prop, f b = if b then True else P. Generating a proof that f is proper, using bool as a setoid relating true and false, and Prop as a setoid with the relation iff, is equivalent to proving iff P True, which is undecidable.
Presently, our automation constructs proper proofs automatically in two practical cases. First, when f is a composition of proper functions, we introduce hypotheses stating that all arguments are equivalent, and then rewrite by these hypotheses. This is the approach taken by Coq’s
solve_proper tactic, though we modify it slightly. Coq’s solve_proper fails if any of its inputs do not appear in the body of the function because the rewrite for that argument will fail. To avoid this, we use the try tactical when rewriting. We also try both Coq’s rewrite and setoid_rewrite tactics, while solve_proper only runs setoid_rewrite. While setoid_rewrite supports rewriting under binders in some instances, rewrite succeeds in some instances where setoid_rewrite fails.
Second, suppose that f is an application of the eliminator of some inductive type with a constant motive. Then, if all of the inductive case arguments provided are proper functions, we can prove that the eliminator is a proper function from its base cases to its output. To make this concrete, we will consider an example. Let (C, eqC) be a setoid, fix some P = fun _ => C, and consider:
nat_rect P : $\forall$ (po : C) (ps : forall (n : nat) (pn : C), C) (n : nat), C
@@
Proper (eq ==> eqC ==> eqC) ps ->
$\forall$ (n : nat), Proper (eqC ==> eqC) (fun po => nat_rect P po ps n)
by induction on n. The base case holds by assumption via the definition of Proper, and the inductive case holds by ps being proper.
This can be done in general for inductive types. Thus, our automation checks if f is an eliminator at the top level, and if so generates and tries to prove such a proper goal. If it succeeds, the automation attempts to show that the inductive cases (in the above example, ps) are proper using the rewriting strategy described above. In our case studies, this is necessary for automatically solving some of the generated proper goals.
@@
Figure 8. The types of our integer representations. We provide an instance of Equivalence eq_GZ for the case study in the supplementary material.
@@ 5. Case Studies
We use our extended version of Pumpkin Pi to automatically repair proofs on three case studies that use quotient type equivalences. First, we conduct repair between two representations of the integers (Section 5.1). Second, we study two common implementations of the queue data structure and how we can repair from one to the other (Section 5.2). Third, we repair between dense and sparse representations of polynomials with natural number coefficients (Section 5.3). In each of these case studies, the type to which we repair functions and theorems has a structure which enables efficient implementations of key operations. Our extension makes it possible to take advantage of this structure while using the repaired theorems, something impossible with prior versions of Pumpkin Pi. All of the case study examples can be found in more detail in the supplementary material.
@@ 5.1. Adding, Fast and Slow
Our first case study is a usage of quotients that is foundational in mathematics. We consider a change in the type representing integers from the inductive type found in many standard libraries to the standard quotient based representation. We repair addition and proofs about addition from one representation to the other. Finally, we recover the repaired proofs for a more efficient version of addition over the repaired type. Our Coq implementation of this case study can be found in grothendieck_int_equivalence_repair_tool.v.
@@
Our first representation, Z, is based on the default implementation of the integers in Cubical Agda: two copies of $\mathbb{N}$ glued back-to-back. We will repair functions and proofs about Z to use a representation that may be more familiar to set theorists: viewing the integers as elements of $\mathbb{N}\times\mathbb{N}/\sim$, where $(x_{1},x_{2})\sim(y_{1},y_{2})\iff x_{1}+y_{2}=x_{2}+y_{1}$. We call the resulting type GZ [^3] In reference to Grothendieck, as this is the Grothendieck group of the natural numbers, and the equivalence relation eq_GZ. The definitions are in Figure 8. Z and GZ are setoid equivalent using the definition from Section 2, by mapping pos n to (n, 0) and negsuc n to (0, S n).
@@
Figure 10. An addition identity whose proof we repaired automatically.
@@
We use appliedDepElimPropZ in our proof of add0LZ, and can repair the term. Presently, an implementation bug only surfacing in this case study forces us to repair the arguments to
appliedDepElimPropZ before reconfiguring. Future versions can avoid the need for this workaround by automatically supplying the necessary proper proofs to depElimProp terms.
@@
We adapt our repaired proofs to use the more efficient addition function defined in Figure 11. Consistently with prior work in Pumpkin Pi, to move between slow and fast implementations we take an ad hoc approach. First, we prove that both implementations of addition are pointwise equivalent in our setoid, producing a term addEqualFastAdd. Then, we rewrite across this equality to get from proofs of repaired theorems defined over slow addition to proofs of corresponding theorems defined over fast addition. For example, we use this methodology to translate the proof of add0LGZ into the proof of theorem:
@@
Figure 12. One list queues and two list queues. We provide an instance of Equivalence eq_queue in the supplementary material.
Definition enqueueOLQ (a : A) (q : OLQ) : OLQ :=
depConstrOLQInsert a q.
Definition dequeueHelpOLQ (outer : A)
(q : OLQ) (m : option (OLQ * A)) :
option (OLQ * A) :=
@option_rect
(OLQ * A)
(fun _ => option (OLQ * A))
(fun (p : (OLQ * A)) => Some
(depConstrOLQInsert outer (fst p) ,
(snd p)))
(Some (depConstrOLQEmpty, outer))
m.
Definition dequeueOLQ :
OLQ -> option (OLQ * A) :=
depRecOLQ (option (OLQ * A)) None dequeueHelpOLQ.
Definition enqueueTLQ (a : A) (q : TLQ) : TLQ :=
depConstrTLQInsert a q.
Definition dequeueHelpTLQ (outer : A)
(q : TLQ) (m : option (TLQ * A)) :
option (TLQ * A) :=
@option_rect
(TLQ * A)
(fun _ => option (TLQ * A))
(fun (p : (TLQ * A)) => Some
(depConstrTLQInsert outer (fst p) ,
(snd p)))
(Some (depConstrTLQEmpty, outer))
m.
Definition dequeueTLQ :
TLQ -> option (TLQ * A) :=
depRecTLQ (option (TLQ * A)) None dequeueHelpTLQ.
Figure 13. Definitions for enqueue and dequeue over OLQ on the left, and their repaired versions over TLQ on the right.
Definition returnOrEnqOLQ (a : A)
(m : option (OLQ * A)) : (OLQ * A) :=
@option_rect
(OLQ * A)
(fun _ => prod OLQ A)
(fun (p : (OLQ * A)) =>
(enqueueOLQ a (fst p), snd p))
(depConstrOLQEmpty, a)
m.
Theorem dequeueEnqueue (a : A) (q : OLQ) :
dequeueOLQ (enqueueOLQ a q)
= Some (returnOrEnqOLQ a (dequeueOLQ q)).
Definition returnOrEnqTLQ : (a : A)
(m : option (TLQ * A)) : (TLQ * A) :=
@option_rect
(TLQ * A)
(fun _ => prod TLQ A)
(fun (p : (TLQ * A)) =>
(enqueueTLQ a (fst p), snd p))
(depConstrTLQEmpty, a)
m.
Theorem dequeueEnqueue (a : A) (q : TLQ) :
eq_deq_ret
(dequeueTLQ (enqueueTLQ a q))
(Some
(returnOrEnqTLQ a (dequeueTLQ q))).
Figure 14. Main theorem relating dequeue and enqueue, stated for OLQ on the left and TLQ on the right. We repair this theorem’s proof from OLQ to TLQ. Here, eq_deq_ret is the equivalence relation eq_queue lifted to the return type of dequeueTLQ.